# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v4

@@ II-B Radar system
The radar system is designed for high adaptability, allowing for real-time adjustments in detection capabilities across $L$ frame snapshots. $T$ represents the radar beams emitted during a specific detection frame. Using MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [33, 34], where $T$ signifies the upper limit of detectable targets in a single frame.
@@
$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vspace{-0.05 in}$ (4)
@@
$\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}\>q+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}\>q+\mathbf{n}_{t,l},$ (5)
@@ III The communication KLD: ZF and Arbitrary precoding matrix
In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver. The analysis is carried out for the normalised precoders introduced in Section II-A. To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
@@ III-A ZF beamforming
With ZF precoding, the received signal at the $k$-th user is,
@@
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix. Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI. The distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
$\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}P_{\mathrm{r}}\sigma_{h}^{2}+\sigma_{n}^{2}$ (8)
where $\mathbf{R}_{w,t}\!\!=\!\!\frac{1}{L}\!\sum_{l=1}^{L}\!\!\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}^{H}$. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
@@
$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (9)
@@
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (9), and (10) in (6), and considering all possible pairs of dissimilar symbols [22, Corollary 1], then the KLD becomes,
@@
[…]
@@ III-B Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem. Considering the received signal representation in (2), the conditional mean vectors and variance are,
$\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\$ (14)
@@
$=\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\!\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2},\!\!\!\quad j\in\{n,m\},$ (15)
Note that, the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
@@
×$\!\!\times\!\Bigg{(}\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg{)}^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg{)}\!.\!\!$ (17)
Thereafter, by considering all possible pairs of dissimilar symbols, then the KLD expression using $\lambda$ is simplified as,
@@
It should be noted that the second term in this equation, represents a form of the SINR, where the interference is the sum of IUI and radar interference.
@@ IV-A Composite detection and response matrix estimation
The main objective of typical radar systems is to detect the existence of targets and estimate the matrix response for existing targets. After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
@@
This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining ^$\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if presents. It can be observed from (5) that the distribution of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis is multivariate normal whose PDF is,
@@
×$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (20)
@@ IV-B The KLD of radar system
[…]
@@ V Radar Waveform Optimisation
In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem is,
$\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\quad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (29a)
@@
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (27). After substituting (27) in (29a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance. The minimum KLD requirements in constraints (29b) and (29c) are carefully chosen to ensure specific performance guarantees in terms of detection probability $P_{D}$ and BER for the radar and communication subsystems, respectively. The constraint on radar power (29d) is explicit because the radar waveform is our primary optimisation variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (29c) on communication performance. Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
@@
The gradient of the objective function, $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$, is a gradient tensor computed element-wise, as detailed in Appendix I.A. The final equation is given by,
@@
where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function. The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A. The final forms are given by,

### v5

@@ II-B Radar system
The radar system is designed for high adaptability, allowing real-time adjustments in detection capabilities across L\!L\! frame snapshots. TT\! denotes the radar beams emitted during a given detection frame. Using MIMO radar technology, multiple beams can be formed simultaneously via orthogonal signals [41, 42], where $T$ is the upper limit of detectable targets in one frame.
@@
$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vskip-3.61371pt$ (3)
@@
$\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\overbrace{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}}^{\begin{subarray}{c}\text{Radar waveform}\\ \text{ radar return}\end{subarray}}\>q+\overbrace{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}}^{\begin{subarray}{c}\text{Comms. waveform}\\ \text{ radar return}\end{subarray}}q+\overbrace{\mathbf{n}_{t,l}}^{\text{Noise}},$ (4)
@@ III The communication KLD: ZF and Arbitrary precoding matrix
In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, motivated by its direct connection to the maximum likelihood detection process and tractable analytical expressions for system optimisation, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver. The analysis is carried out for the normalised precoders introduced in Section II-A. To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
@@ III-A ZF beamforming
ZF designs are typically used to precode the communication data symbols at the BS. This scheme will serve as a benchmark for our designs. ZF eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This usually results in excellent performance at high SNR values. The normalised ZF beamforming matrix is also designed based on the channel matrix, and can be represented as follows,
where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE. With ZF precoding, the received signal at the $k$-th user is,
@@
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of $\mathbf{W}_{\mathrm{c}}$. It is worth noting that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI. The radar interference term $\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}$ consists of $NT$ independent random variables arising from the random radar precoding. For moderate to large $NT$ values, CLT ensures that this sum converges to a Gaussian distribution. Combined with the inherently Gaussian noise $n_{k,l}$, the total interference-plus-noise $\eta_{k}$ is well-approximated as $\mathcal{CN}(0,\sigma_{\eta}^{2})$, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of
$\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}\sigma_{h}^{2}\mathrm{Tr}(\mathbf{R}_{w})+\sigma_{n}^{2}$ (8)
The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
@@
$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\qquad j\in\{n,m\},$ (9)
@@
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter NK+1N-K+1 and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (9), and (10) in (5), and considering all possible pairs of dissimilar symbols [23, Corollary 1], then the KLD becomes,
@@
$\begin{split}\mathbb{E}\!\left[\|\mathbf{\tilde{W}}_{\!\mathrm{c}}\|_{F}^{-2}\right]\!\!&=\!\!\!\int_{0}^{\infty}\!\!\frac{1}{x}\cdot\frac{1}{\Gamma(L_{\mathrm{G}})}x^{L_{\mathrm{G}}-1}e^{-x}\!dx\!=\!N\!\!-\!\!K.\end{split}$ (12)
@@ III-B Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem. By noting that the received signal representation $y_{\mathrm{c},k,l}$ in (2) consists of a desired term, (K1)(K-1) statistically independent inter-user symbols and $NT$ independent radar/noise terms, as well as, by considering CLT, the resulting sum is well approximated by a proper complex Gaussian for a typical MIMO-ISAC system which employs large antenna arrays. We therefore take $y_{\mathrm{c},k,l}^{(j)}\sim\mathcal{CN}(\mu_{k,j},\Sigma_{\mathrm{c},j}),\ \ j\in\{n,m\}$, where the conditional mean vectors and variance are,
$\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\\$ (14)
@@
$=\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\!\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2},\!\!\!\qquad j\in\{n,m\}.$ (15)
Note the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
@@
$\!\!\times\!\Bigg(\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg)^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg)\!.\!\!$ (17)
Thereafter, by considering all possible pairs of dissimilar symbols, the KLD expression using $\lambda$ is simplified as,
@@
Where the interference is the sum of IUI and radar interference. It should be noted that the second term in this equation represents a form of the SINR, confirming KLD’s alignment with conventional metrics while uniquely enabling modulation-aware distinguishability and synergistic interference exploitation through its unified hypothesis-testing foundation.
@@ IV-A Composite detection and response matrix estimation
The main objective of radar systems is to detect the existence of targets and estimate the matrix response for existing targets. After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
@@
This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining $\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if present. It can be observed from (4) that the distribution of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis is multivariate normal whose PDF is,
@@
$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (20)
@@ IV-B The KLD of radar system
As the number of snapshots $L$ approaches infinity, the law of large numbers ensures that the estimated target response matrix converges in probability to the true value, i.e., $\hat{\mathbf{H}}_{t}\xrightarrow{p}\mathbf{H}_{t}$. This allows us to use $\mathbf{H}_{t}$ in place of $\hat{\mathbf{H}}_{t}$ for large $L$. The received radar signal $\mathbf{y}_{r,t,l}$ becomes a linear transformation of the Gaussian noise $\mathbf{n}_{t,l}$ plus deterministic signal components, where we can proceed with a more tractable derivation of the KLD for the radar subsystem.
@@
This formulation represents the aggregate information-theoretic performance measure across both subsystems, where the averaging within each subsystem ensures balanced treatment regardless of the number of users $K$ or targets $T$, where $\mathrm{KLD}_{\mathrm{ISAC}}$ provides a unified optimisation objective that captures the fundamental goal of maximising information extraction from shared ISAC resources.
@@ V Radar Waveform Optimisation
The KLD-based optimisation framework provides analytically tractable KLD expressions, and their gradients as shown in Sec. III, Sec. IV-B, and Appendix A. This computational efficiency, combined with the unified performance characterisation, makes KLD particularly suitable for real-time ISAC applications. In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem is,
$\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\qquad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (29a)
@@
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (27). After substituting (27) in (29a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance. The minimum KLD requirements in constraints (29b) and (29c) are carefully chosen to ensure specific performance guarantees in terms of detection probability $P_{D}$ and BER for the radar and communication subsystems, respectively. For example, for the radar subsystem, a KLD threshold of $\underline{A}_{t}=10$ bits corresponds to a detection probability of approximately $P_{D}\approx 0.25$, representing a baseline detection capability. For the communication subsystem, a KLD threshold of $\underline{B}_{k}=10$ bits ensures a BER of approximately $10^{-2}$ for QPSK modulation, which is acceptable for many practical applications. These thresholds can be adjusted based on specific system requirements—higher KLD values would yield improved performance at the cost of more stringent resource allocation. The mathematical relationships between KLD and these conventional metrics, established in [23], allow system designers to set appropriate thresholds based on application needs. The constraint on radar power (29d) is explicit, where fixed power allocation is used between subsystems, because the radar waveform is our primary optimisation variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (29c) on communication performance. Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
@@
The gradient of the objective function, $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$, is a gradient tensor computed element-wise, as detailed in Appendix A The final equation is given by,
@@
where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function. The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix A. The final forms are given by,
@@
V-A Radar waveform detection probability optimisation
To incorporate the detection performance directly into the optimisation framework, we derive the probability of detection $P_{D,t}$ from the GLRT in Sec. IV.A, where $\mathbf{Q}_{t}=\mathbf{I}_{N}+\mathbf{R}_{1,t}^{-1}$. For large $L$ and by invoking the CLT, the test statistic converges to a Gaussian distribution under both hypotheses. Under $\mathcal{H}_{0}$ where $\mathbf{y}_{\mathrm{r},t,l}=\mathbf{n}_{t,l}$, the mean and variance are $\mu_{0}=\sigma_{n}^{2}\text{tr}(\mathbf{Q}_{t})$ and $\sigma_{0}^{2}=\sigma_{n}^{4}\text{tr}(\mathbf{Q}_{t}^{2})/L$. Under $\mathcal{H}_{1}$ where $\mathbf{y}_{\mathrm{r},t,l}=d_{\mathrm{r},t}^{\zeta/2}\mathbf{H}_{t}\mathbf{x}_{t,l}+\mathbf{n}_{t,l}$, we have $\mu_{1}=\mu_{0}+\Delta\mu$ and $\sigma_{1}^{2}=\sigma_{0}^{2}+2\sigma_{n}^{2}\Delta\mu/L$, where $\Delta\mu=d_{\mathrm{r},t}^{\zeta}\text{tr}(\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}\mathbf{Q}_{t})$. For a fixed false alarm rate $P_{FA}$, the Neyman-Pearson threshold is,
$\tau_{2}=\mu_{0}+Q^{-1}(P_{FA})\sigma_{0}.$ (42)
Substituting (42) into the detection probability expression $P_{D,t}=Q((\tau_{2}-\mu_{1})/\sigma_{1})$ yields [23],
$P_{D,t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}=Q\left(\frac{Q^{-1}(P_{FA})\sigma_{0}-\Delta\mu}{\sigma_{1}}\right),$ (43)
where $Q(\cdot)$ is the complementary cumulative distribution function of the standard normal distribution. The detection probability in (43) depends on the waveform design through $\mathbf{R}_{t}$, which affects both $\mathbf{Q}_{t}$ and $\Delta\mu$. We form a baseline radar waveform optimisation by using the detection probability directly in the problem defined in (44) below.
$\mathcal{P}_{3}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\qquad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}P_{D,t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (44a)
[…]
$P_{D,t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}\leq\overline{D}_{t}\,\forall t\in T,$ (44b)
where the communication subsystem uses ZF beamforming with precoding matrix $\mathbf{W}_{\mathrm{c}}$ and $\overline{D}_{t}$ is the upper limit of the $P_{D}$ of the $t$-th target. This $P_{D,t}$ objective depends on the $Q(\cdot)$ function and its inverse, which introduces additional computational complexity in the optimisation process due to its non-elementary nature and the challenges in evaluating its derivatives efficiently. This distinction illustrates the enhanced analytical tractability of the proposed KLD-based approach. This problem is inherently non-convex and nonlinear due to the $Q$-function and its constraint. Using $\nabla_{\mathbf{W}_{\mathrm{r},t}}\,\mathrm{P}_{D,t}$ derived in Appendix B, and the gradients derived previously, we employ the same projected gradient method with penalty as in Algorithm 1, to handle the nonconvexities in (44).